OpenStudy (anonymous):

If f(x) = x + 2, find f(3). 1 6 5 −1

OpenStudy (anonymous):

Somebody please help me on this!!!

OpenStudy (anonymous):

The answer is 5. f(x) = x + 2 f(3) = 3 + 2 f(3) = 5

OpenStudy (akashdeepdeb):

f(x) = x+2 f(a) = a+2 f(b) = b+2 f(3) = 3+2 = 5 Just put the value of x in the function too. ^_^ Understood the concept? :D

OpenStudy (anonymous):

Yes thank you(:

OpenStudy (akashdeepdeb):

:)